# FlipedRAG: Black-Box Opinion Manipulation Attacks to Retrieval-Augmented Generation of Large Language Models

### v1

@@
Abstract
Retrieval-Augmented Generation (RAG) addresses hallucination and real-time constraints by dynamically retrieving relevant information from a knowledge database to supplement the LLMs’ input. When presented with a query, RAG selects the most semantically similar texts from its knowledge bases and uses them as context for the LLMs to generate more accurate responses. While RAG offers numerous benefits, it also creates a new attack surface for large language models, especially since RAG databases are frequently sourced from public domains such as the web. While existing studies have predominantly focused on optimizing RAG’s performance and efficiency, emerging research has begun addressing the security concerns associated with RAG. However, these works have some limitations, typically focusing on either white-box methodologies or heuristic-based black-box attacks, and often lacking of in-depth exploration of retrieval subsystem. Furthermore, prior research has mainly targeted simple factoid question answering, which is neither practically challenging nor resistant to correction by LLMs. In this paper, we unveil a more realistic and threatening scenario: opinion manipulation for controversial topics against RAG. Particularly, we propose a novel RAG black-box attack method, termed FlipedRAG, which is transfer-based. By leveraging instruction engineering, we obtain partial retrieval model outputs from black-box RAG system, facilitating the training of surrogate models to enhance the effectiveness of opinion manipulation attack. Extensive experimental results confirms that, compared to the baseline method, our approach significantly enhances the average success rate of opinion manipulation by 16.7%. It achieves an average of a 50% directional change in the opinion polarity of RAG responses across four themes. Additionally, it induces a 20% shift in user cognition. Furthermore, we discuss the efficacy of potential defense mechanisms and conclude that they are insufficient in mitigating this type of attack, highlighting the urgent need to develop novel defensive strategies.
1 Introduction
Figure 1: An overview of our black-box opinion manipulation for revealing the vulnerabilities of RAG and its potential impact on people’s cognition.
With the rapid development of artificial intelligence, large language models (LLMs) have demonstrated exceptional capabilities in the field of natural language processing. However, constrained by their training data, these models have limited scope of knowledge and lack the most up-to-date information, which can lead to errors or hallucinations when tackling more complex or time-sensitive tasks. Retrieval-Augmented Generation (RAG) combines information retrieval with the generative capabilities of LLMs, enhancing the timeliness of knowledge acquisition and effectively mitigating the hallucination problem of these models. When given a query, RAG retrieves the most relevant documents from a knowledge base to augment the input request for the LLM. For example, the retrieved knowledge may consist of a series of text snippets that are semantically most similar to the query. RAG has inspired many popular applications, such as Microsoft Bing Chat, ERNIE Bot, and KimiChat, which use RAG to summarize retrieval results for improving user experience. Open-source projects like LangChain and LlamaIndex provide developers with flexible RAG frameworks to build customized AI applications using LLMs, retrieval models and knowledge bases.
@@ 1 Introduction
As the application scope of RAG expands, concerns about its security, particularly regarding model performance under malicious attacks, have become prominent. The fundamental RAG process generally consists of three components: knowledge bases (corpus collections), retriever, and LLMs. Vulnerabilities arise when some of the retrieved documents are compromised by malicious attackers, a scenario referred to as a manipulation attack against RAG in this paper. Numerous studies have explored various forms of adversarial manipulation attacks, such as adversarial attack on the retriever [13, 15], prompt injection attack [1, 14, 9], jailbreak attack for LLM [4, 11, 26], and poisoning attack targeting the retrieval corpus in RAG [29, 23].
This paper primarily examines adversarial ranking poisoning attacks targeting the retriever in black-box RAG systems, which align with realistic and practical scenarios. It explores how such attacks indirectly influence the opinionated generative outputs of LLMs, as illustrated in Figure 1. By employing black-box retriever imitation [13] for opinion manipulation, we reveal specific vulnerabilities in RAG systems. The threat model presented here closely mirrors a real-world black-box scenario and can be characterized as follows: the adversary can only query the RAG system and does not have access to the complete knowledge base or corpus, the retriever, or the parameters of the RAG. The attacker is only capable of injecting adversarially modified candidate texts into the corpus, while the retriever and the LLM remain black-boxed, intact, and unmodifiable. Based on previous research [12, 2], the retrieval corpus and knowledge base consist of millions of candidate texts derived from the internet, enabling attackers to insert adversarially modified content through the malicious crafting of web content or encyclopedic entries. Cho et al. [3] and Zhong et al. [27] investigate the attack against predefined retrievers in white-box RAG, which are challenging to achieve in real-world scenarios with limited flexibility and practicality. Furthermore, they did not evaluate attacks specifically targeting the integrated generation process. Another representative work, PoisonedRAG [29], implemented black-box retrieval poisoning attacks against RAG knowledge bases, effectively uncovering pertinent security vulnerabilities. However, it predominantly concentrated on closed-domain question answering, such as "Who is the CEO of OpenAI?" These types of questions could be rectified when RAG systems are coupled with fact-checking and value alignment in LLMs.
@@
The vulnerabilities explored in this paper primarily address controversial and opinion-based inquiries in RAG, such as "Should abortion be legal?" These questions are open-ended and require advanced levels of logical reasoning and summarization capabilities from LLMs. Current research on controversial topics is limited, and attacks that manipulate opinions on such questions could potentially inflict deeper harm.
Controversial topics are issues characterized by a lack of consensus due to divergent perspectives, often attracting widespread attention. These topics frequently involve differing viewpoints that shape public perception during extensive discussions. For instance, in political elections, Robert Epstein [6] demonstrated that manipulating search engines to generate biased results can alter voters’ preferences. In scenarios involving controversial topics, opinion manipulation can create “Information Bubbles”, which subsequently lead to the homogenization of user opinions [24]. In such cases, individuals’ views become easily swayed by the stance of the information they are exposed to. With the increasing use of LLMs in content production, exploiting vulnerabilities in RAG systems for opinion manipulation represents a particularly grave threat. Through carefully crafted inputs, attackers can influence the orientation of content generated by these models, thus jeopardizing users’ cognitive processes and decision-making abilities. Consequently, it is of considerable theoretical and practical importance to examine the vulnerabilities of RAG models to opinion manipulation attacks within a black-box setting.
@@
To address aforementioned challenges, in this paper, we propose FlipedRAG, a black-box attack method, to explore the reliability of RAG in controversial topics and investigate its impact on user cognition. Specifically, we begin by inputting carefully designed instructions to obtain the rankings of retrieval results within the RAG model, thereby analyzing the operational mechanism of its retrieval module. Subsequently, we train a surrogate model using the extracted retrieval data [13, 20] to approximate the features and relevance preferences of the retriever in the black-box RAG system. Based on this surrogate model, we develop an adversarial attack strategy aimed at manipulating the opinions of candidate documents. By attacking this surrogate model, we generate adversarial opinion manipulation triggers, which are then transferred to the target RAG corpus. We conduct experiments on opinion datasets encompassing multiple topics to validate the effectiveness and scope of the attack strategy without relying on internal knowledge of the RAG model. The experimental results reveal that the proposed attack strategy can significantly alter the opinion polarity of the content generated by RAG. Furthermore, we discuss the efficacy of potential defense mechanisms and conclude that they are insufficient in mitigating this type of attack. This not only highlights the vulnerability of the retrieval model and the retrieval-augmented generation process but, more importantly, underscores the potential negative impact on user cognition and decision-making. Such vulnerabilities can increase the risk of misleading users into accepting incorrect or biased information.
@@
(1) We propose a novel transfer-based opinion manipulation attack against RAG of LLMs under black-box setting, termed FlipedRAG, which aligns more closely with real-world and practical scenarios.
(2) To the best of our knowledge, we are the first to explore the adversarial opinion manipulation of RAG on open-ended and controversial topics.
(3) Through both automated evaluations and user experiments, we demonstrate that our transfer-based manipulation attack can significantly influence the opinions generated by RAG system in a specific desired direction.
(4) We discuss the efficacy of potential defense mechanisms and conclude that they are insufficient in mitigating this type of attack, highlighting the need for new defense strategy.
Figure 2: The overview of FlipedRAG for manipulating the opinions of RAG-generated content under black-box setting.
2 Related Works
2.1 Reliability of Neural Network Models
@@
Research on the reliability of neural network models has been extensively studied. In 2013, Szegedy et al. [17] found that applying imperceptible perturbations to a neural network model during a classification task was sufficient to cause classification errors in CV. Later, scholars observed similar phenomena in NLP. Robin et al. [10] found that inserting perturbed text into original paragraphs significantly distracts computer systems without altering the correct answer or misleading humans. It reflects the robustness of neural network models, i.e., the ability to output stable and correct predictions in tackling the imperceptible additive noises [19]. For large language models, Wang et al. [18] proposed a comprehensive trustworthiness evaluation framework for LLMs, assessing their reliability from various perspectives such as toxicity, adversarial robustness, stereotype bias, and fairness. While large language models have greater capabilities compared with general deep neural network models, they also raise more concerns regarding security and reliability.
2.2 Reliability of RAG
As RAG is designed to overcome the hallucination problem in LLMs and enhance their generative capabilities, the reliability of content generated by RAG also becomes a major concern. Zhang et al. [25] sought to identify the weaknesses of RAG by analyzing critical components in order to facilitate the injection of the attack sequence and crafting the malicious documents with a gradient-guided token mutation technique. Xiang et al. [21] proposed an isolate-then-aggregate strategy, which first obtains responses from each passage of the LLM in isolation and then securely aggregates these isolated responses, to construct the first defense framework against retrieval corruption attacks. These studies are based on white-box scenarios and primarily focus on the robustness of RAG against corrupted and toxic content.
This paper proposes the use of adversarial retrieval attack strategies to manipulate the ranking results of the retriever, ensuring that opinion documents with a specific stance are ranked as highly as possible, thereby influencing the generated responses of the LLM to align with that stance.
2.3 Adversarial Retrieval Attack Strategy
@@
The adversarial retrieval attack strategy starts with manipulation at the word level. Under white-box setting, Ebrahimi et al. [5] utilize an atomic flip operation, which swaps one token for another, to generate adversarial text(known as adversarial trigger) and the method is called Hotflip. Hotflip eliminates reliance on rules, but the adversarial trigger it generates usually has incomplete semantics and insufficient grammar fluency. While it can deceive the target model, it cannot evade perplexity-based defenses. Wu et al. [20] also proposed a word substitution ranking attack method called PRADA. To enhance the readability and effectiveness of the trigger, scholars further designed sentence-level ranking attack methods. Song et al. [16] propose an adversarial method under white-box setting, named Collision, which uses gradient optimization and beam search to produce the adversarial trigger named collision. The Collision method further imposes a soft constraint on collision generation by integrating a language model, reducing the perplexity of the collision. The method has shown promising results in document retrieval experiments. Inspired by Collisions, Liu et al. [13] propose the Pairwise Anchor-based Trigger (PAT) method under black-box setting. Added the fluency constraint and the next sentence prediction constraint, the method generates adversarial triggers by optimizing the pairwise loss of top candidates and target candidates with triggers. Although the time complexity of PAT has increased compared to previous methods, PAT takes ranking similarity and semantic consistency into account, so its manipulation effect on the retrieval ranking of target candidates is superior.
3 Method
3.1 Overview
@@
Our objective is to manipulate the opinions expressed in responses generated by black-box RAG models on controversial topics, with a focus on both the retrieval model and the large language model that performs the integrated generation task. Zhang et al. [25] attempted to poison context documents to mislead the LLM into generating incorrect content. However, this approach necessitates extensive internal details of the LLM application, rendering it less feasible in real-world scenarios. In the black-box RAG context, the attacker lacks access to internal information of the RAG, including model architecture and scoring functions, and can only interact with the inputs and outputs of the RAG. Specifically, the attacker can only use the interface of the RAG and not directly access the retriever. Since the inputs to LLMs are comprised of the user query and context documents, and the user query is immutable, our attack strategy focuses on modifying candidate documents within the corpus. Although the attacker does not have access to the entire corpus, they can insert adversarially modified candidate texts into it, as many RAG applications source information from the Internet, where the content is publicly editable.
@@
The basic framework of RAG comprises two components: the retriever and the LLM. These modules are serially connected, with the retriever sourcing context information from a knowledge base, upon which the LLM then performs the generation task. In black-box scenarios, attackers cannot modify the system prompts of the generative LLM, making it challenging to directly influence the generation results by exploiting any reliability flaws within the LLM itself. Therefore, we focus on exploiting the reliability flaws of the retriever to manipulate the retrieval ranking results. We employ specific instructions to induce RAG to reveal the context information it references. With this information, we train a surrogate model that approximates the black-box retriever, effectively transparentize it into a white-box model. We then employ this white-box model to generate adversarial triggers. By adding these adversarial triggers to candidate documents that reflect the desired opinion, we enhance their relevance to the user query, increasing the likelihood that they will be included in the context passed to the generative LLM. Leveraging the strong capability of LLM for context understanding and instructions-following, we guide it to generate responses that align with the expected opinion. An overview of this process is shown in Figure 2.
3.2 Threat Model

### v3

# FlippedRAG: Black-Box Opinion Manipulation Adversarial Attacks to Retrieval-Augmented Generation Models

@@
Abstract.
Retrieval-Augmented Generation (RAG) enriches LLMs by dynamically retrieving external knowledge, reducing hallucinations and satisfying real-time information needs. While existing research mainly targets RAG’s performance and efficiency, emerging studies highlight critical security concerns. Yet, current adversarial approaches remain limited, mostly addressing white-box scenarios or heuristic black-box attacks without fully investigating vulnerabilities in the retrieval phase. Additionally, prior works mainly focus on factoid Q&A tasks, their attacks lack complexity and can be easily corrected by advanced LLMs. In this paper, we investigate a more realistic and critical threat scenario: adversarial attacks intended for opinion manipulation against black-box RAG models, particularly on controversial topics. Specifically, we propose FlippedRAG, a transfer-based adversarial attack against black-box RAG systems. We first demonstrate that the underlying retriever of a black-box RAG system can be reverse-engineered and approximated by enumerating critical queries, candidates, and answers, enabling us to train a surrogate retriever. Leveraging the surrogate retriever, we further craft target poisoning triggers, altering vary few documents to effectively manipulate both retrieval and subsequent generation, transferring the attack to the original black-box RAG model. Extensive empirical results show that FlippedRAG substantially outperforms baseline methods, improving the average attack success rate by 16.7%. Across four diverse domains, FlippedRAG achieves on average a 50% directional shift in the opinion polarity of RAG-generated responses, ultimately causing a notable 20% shift in user cognition. Furthermore, we actively evaluate the performance of several potential defensive measures, concluding that existing mitigation strategies remain insufficient against such sophisticated manipulation attacks. These results highlight an urgent need for developing innovative defensive solutions to ensure the security and trustworthiness of RAG systems.
RAG, Black-box Adversarial Attack, Opinion Manipulation
*Corresponding author
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacy Software and application securityccs: Information systems Information retrievalccs: Computing methodologies Natural language processing
1. Introduction
Retrieval-Augmented Generation (RAG) combines information retrieval with the generative capabilities of LLMs, enhancing the timeliness of knowledge acquisition and effectively mitigating the hallucination problem of these models (Gao et al., 2023; Zhao et al., 2024b).
@@ 1. Introduction
As the application scope of RAG expands, concerns about its security are also increasing. The basic RAG process typically consists of three components: a knowledge base (corpus collection), a retriever, and an LLM. Vulnerabilities arise when some of the retrieved documents are compromised, or prompt are tampered by malicious attackers, a scenario referred to as a manipulation attack against RAG. Previous studies have explored various forms of adversarial manipulation attacks, such as adversarial attack on the retriever (Liu et al., 2023b, c), prompt injection attack (Cai et al., 2022; Liu et al., 2023a; Jain et al., 2023), jailbreak attack for LLM (Deng et al., 2023; Li et al., 2023; Zhao et al., 2024a), and poisoning attack targeting the retrieval corpus in RAG (Cho et al., 2024; Zou et al., 2024; Xue et al., 2024).
Figure 1. (a) ChatGPT can be induced through user prompts to fully replicate the context segments referenced in its responses. Based on this observation, we can transparentize the retriever of the black-box RAG model by enumerating critical queries and candidates. (b) An overview of our black-box opinion manipulation adversarial attacks to RAG models and its potential impact on user’s cognition.
However, previous studies (Zou et al., 2024; Xue et al., 2024; Cho et al., 2024) have primarily focused on attacks against RAG systems, which lack feasibility and have fundamental limitations. They mostly address white-box scenario or employ heuristic-based black-box attacks, without thoroughly analyzing the vulnerabilities within the retrieval stage. Furthermore, earlier research has mainly concentrated on factoid question answering tasks, such as questions like “Who is the CEO of OpenAI?” Those factual queries are less impactful, as RAG systems combined with fact-checking and value alignment mechanisms in LLMs are capable of rectifying these errors (Xiang et al., 2024; Tu et al., 2025).
@@
Differently, the scenario explored in this paper primarily addresses controversial and opinion-based inquiries in RAG, such as “Should abortion be legal?” These questions are open-ended and require advanced levels of logical reasoning and summarization capabilities from LLMs. Current research on controversial topics is limited. In scenarios involving controversial topics, opinion manipulation can create “Information Bubbles”, which subsequently lead to the homogenization of user opinions (Yue et al., 2023). In such cases, individuals’ views become easily swayed by the stance of the information they are exposed to. Attacks that manipulate opinions on such questions could potentially inflict deeper harm (Epstein and Robertson, 2015). Through carefully crafted inputs, attackers can influence the orientation of content generated by RAG models, thus jeopardizing users’ cognitive processes and decision-making abilities.
Thus, this paper primarily investigates adversarial opinion manipulation targeting the retriever in black-box RAG systems, which align with realistic and practical scenarios. The threat model presented here can be characterized as follows: the adversary can only query the RAG system and does not have access to the complete knowledge base or corpus, the retriever, or the parameters of the RAG. The attacker is only capable of injecting limited adversarially modified candidate texts into the corpus, while the retriever and the LLM remain black-boxed, intact, and unmodifiable.
To address the aforementioned challenges, in this paper, we propose FlippedRAG, a black-box attack method, to explore the reliability of RAG in controversial topics and investigate its impact on user cognition, as shown in Figure 1. By employing black-box retriever imitation (Liu et al., 2023b) for opinion manipulation, we reveal specific vulnerabilities in RAG systems. Our motivation is based on the observation in Figure 1(a). When we queried the RAG on a specific topic and included in the user prompt an instruction requesting LLMs to output the original content of its referenced retrieval results, it would faithfully reproduce the exact text segments it referenced. The reproduced outputs precisely matched the corresponding content on the web pages linked in the reference citations and reflect the functionality of relevance matching. It allows us to design a retriever imitation method for black-box RAG to address the limitations of white-box attacks or heuristic-based attacks against such systems. Specifically, we utilize carefully designed instructions to obtain the retrieval results within the RAG model. Subsequently, based on the constructed pairs (query, candidate) from the retrieval list, we train a surrogate model using the extracted retrieval data (Liu et al., 2023b; Wu et al., 2022b) to approximate and transparentize the relevance preferences of the retriever in the black-box RAG system.
@@
Based on this surrogate model, we develop an attack strategy aimed at manipulating the opinions of candidate documents. By attacking this surrogate model, we generate adversarial opinion manipulation triggers, which are then transferred to the target RAG corpus, as shown in Figure 1 (b). We conduct experiments on opinion datasets encompassing multiple topics to validate the effectiveness and scope of the attack strategy without relying on internal knowledge of the RAG model. The experimental results reveal that the proposed attack strategy can significantly alter the opinion polarity of the content generated by RAG. Furthermore, we discuss the efficacy of potential defense mechanisms and conclude that they are insufficient in mitigating this type of attack. This not only highlights the vulnerability of the retrieval model and RAG model but, more importantly, underscores the potential negative impact on user cognition and decision-making. Such vulnerabilities can increase the risk of misleading users into accepting incorrect or biased information.
@@
(1) We propose FlippedRAG, a novel transfer-based opinion manipulation attack against black-box RAG systems. It realistically reflects practical threat scenarios, and to our knowledge, we are among the first to study adversarial opinion manipulation targeting open-ended and controversial topics in RAG-based models.
(2) We propose a new approach to extract context documents and constructing pseudo-relevant training data for surrogate retriever models in black-box settings. It highlights risks regarding context data leakage in LLM applications. Extensive automated evaluations combined with user studies confirm that our method effectively manipulates the opinions generated by RAG models, ultimately influencing user perception toward specifically targeted viewpoints.
(3) We thoroughly evaluate existing defense measures and find that they fail to effectively prevent or reduce the impact of FlippedRAG, underscoring the urgent need for novel and more robust mitigation strategies.
2. Related Works
2.1. Reliability of Neural Network Models
@@
Research on the reliability of neural network models has been extensively studied. In 2013, Szegedy et al. (Szegedy et al., 2013) found that applying imperceptible perturbations to a neural network model during a classification task was sufficient to cause classification errors in CV. Later, scholars observed similar phenomena in NLP. Robin et al. (Jia and Liang, 2017) found that inserting perturbed text into original paragraphs significantly distracts computer systems without altering the correct answer or misleading humans. It reflects the robustness of neural network models, i.e., the ability to output stable and correct predictions in tackling the imperceptible additive noises (Wang et al., 2019). For LLMs, Wang et al. (Wang et al., 2023) proposed a comprehensive trustworthiness evaluation framework, assessing their reliability from various perspectives such as toxicity, adversarial robustness, stereotype bias, and fairness. While LLMs have greater capabilities compared with general deep neural network models, they also raise more concerns regarding security and reliability.
2.2. Adversarial Retrieval Attack Strategy
@@
The adversarial retrieval attack strategy starts with manipulation at the word level. Under white-box setting, Ebrahimi et al. (Ebrahimi et al., 2017) utilize an atomic flip operation, which swaps one token for another, to generate adversarial text, named Hotflip. Hotflip eliminates reliance on rules, but the adversarial trigger it generates usually has incomplete semantics and insufficient grammar fluency. While it can deceive the target model, it cannot evade perplexity-based defenses. Wu et al. (Wu et al., 2022b) also proposed a word substitution ranking attack method called PRADA. To enhance the readability and effectiveness of the trigger, Song et al. (Song et al., 2020) propose an adversarial method under white-box setting, named Collision, which uses gradient optimization and beam search to produce the adversarial trigger. Collision method further imposes a soft constraint on collision generation by integrating a language model, reducing the perplexity of the collision. Futhermore, Liu et al. (Liu et al., 2023b) propose the Pairwise Anchor-based Trigger (PAT) generation under black-box setting. By adding the fluency and semantic consistency constraints, it generates adversarial triggers by optimizing the pairwise loss of top candidates and target candidates with triggers.
2.3. Reliability of RAG
Some studies have investigated the adversarial poisoning attacks on RAG models, including PoisonedRAG(Zou et al., 2024), BadRAG(Xue et al., 2024), Phantom(Chaudhari et al., 2024), and GARAG(Cho et al., 2024). The distinctive characteristics of these attack methodologies are summarized in the Table 1. “LLM Capability Utilization” refers to the motivation of leveraging LLMs’ capabilities to induce malicious outputs. “Depth” denotes the cognitive penetration depth of RAG attacks. Although both BadRAG(Xue et al., 2024) and Phantom(Chaudhari et al., 2024) target user perception, their manipulations remain confined to sentiment level without achieving deeper ideological persuasion.
As RAG is designed to overcome the hallucination problem in LLMs and enhance their generative capabilities, the reliability of content generated by RAG also becomes a major concern. Zhang et al. (Zhang et al., 2024) sought to identify the weaknesses of RAG by analyzing critical components in order to facilitate the injection of the attack sequence and crafting the malicious documents with a gradient-guided token mutation technique. Xiang et al. (Xiang et al., 2024) proposed an isolate-then-aggregate strategy, named RobustRAG, which first obtains responses from each passage of the LLM in isolation and then securely aggregates these isolated responses, to construct the first defense framework against retrieval corruption attacks. These studies are based on white-box scenarios and primarily focus on the robustness of RAG against corrupted and toxic content.
Table 1. Comparison of existing RAG attacks. DoS denotes denial-of-service.
Methods Query type Scenario LLM Capability Utilization Depth
PoisonedRAG (Zou et al., 2024) Closed-ended, Factoid White-box/Black-box Logical reasoning
BadRAG (Xue et al., 2024) Open-ended, Factoid, DoS White-box Logical reasoning Sentiment
Phantom (Chaudhari et al., 2024) Open-ended, Factoid, DoS White-box Instruction-following Sentiment
GARAG (Cho et al., 2024) Closed-ended, Factoid Gray-box Noise resistence
FlippedRAG Controversial, Open-ended Black-box Induction, logical reasoning Perspectives
This paper proposes the use of adversarial retrieval attack to manipulate the retrieval process in black-box RAG, ensuring that corrupted documents with a specific opinion are ranked as highly as possible, thereby inducing the LLM responses to align with that opinion on controversial topics.
3. Methodology
3.1. Overview
@@
Our objective is to manipulate the opinions expressed in responses generated by black-box RAG models on controversial topics, with a focus on both the retrieval model and the LLM that performs the integrated generation task. Zhang et al. (Zhang et al., 2024) attempted to poison context documents to mislead the LLM into generating incorrect content. However, this approach necessitates extensive internal details of the LLM application, rendering it less feasible in real-world scenarios. In the black-box RAG context, the attacker lacks access to internal information of the RAG, including model architecture and scoring functions, and can only interact with the inputs and outputs of the RAG. Specifically, the attacker can only use the interface of the RAG and not directly access the retriever. Since the inputs to LLMs are comprised of the user query and context documents, and the user query is immutable, our attack strategy focuses on modifying candidate documents within the corpus. Although the attacker does not have access to the entire corpus, they can insert adversarially modified candidate texts into it, as many RAG applications source information from the Internet, where the content is publicly editable.
@@
The basic framework of RAG comprises two components: the retriever and the LLM. These modules are serially connected, with the retriever sourcing context information from a knowledge base, upon which the LLM then performs the generation task. In black-box scenarios, attackers cannot modify the system prompts of the generative LLM, making it challenging to directly influence the generation results by exploiting any reliability flaws within the LLM itself. Therefore, we focus on exploiting the reliability flaws of the retriever to manipulate the retrieval ranking results. We employ specific instructions to induce RAG to reveal the context information it references. With this information, we train a surrogate model that approximates the black-box retriever, effectively transparentize it into a white-box model. We then employ this white-box model to generate adversarial triggers. By adding these adversarial triggers to candidate documents that reflect the target opinion, we enhance their relevance to the user query, increasing the likelihood that they will be included in the context passed to the generative LLM. Leveraging the strong capability of LLM for context understanding, summarization and instructions-following, we guide it to generate responses that align with the expected opinion. An overview of this process is shown in Figure 2. The proposed method is outlined in Algorithm 1 in Appendix A.1.
Figure 2. The overview of FlippedRAG for manipulating the opinions of RAG-generated content under black-box setting.
[…]
FlippedRAG Attack Method Overview.
3.2. Threat Model